OpenStudy (anonymous):

limit as x -->2 (x^3 - 8)/(x^2 - 4)

OpenStudy (anonymous):

factor and cancel

OpenStudy (anonymous):

a^3-b^3=(a-b)(a^2+ab+b^2)

OpenStudy (lgbasallote):

3

OpenStudy (anonymous):

I used L'Hospitals rule and got 3 too

OpenStudy (anonymous):

Factor the numerator and denominator to get (x-2)(x^2+2x+4)/(x-2)(x+2); The x-2 expressions cancel to get (x^2+2x+4)/(x+2) Now use direct substitution to get (2^2+2(2)+4)/(2+2) =3

OpenStudy (anonymous):

Got it, thanks!